# Accelerating Equity: Overcoming the Gender Gap in VC Funding

### v1

@@ Abstract
We examine the growing gender gap in venture capital funding, focusing on accelerator programs in the U.S. We collect a unique dataset with detailed information on accelerators and startups. Using a two-stage methodology, we first estimate a matching model between startups and accelerators, and then use its output to analyze the gender gap in post-graduation outcomes through a control function approach. Our results show that female-founded startups face a significant funding disadvantage, primarily due to relocation challenges tied to family obligations. However, larger cohorts and higher-quality accelerators help reduce this gap by offering female founders better networking opportunities and mentorship.
@@ 1 Introduction
Notes: Using data from Pitchbook, (2021, 2022), we compute VC deal counts and average deal sizes by gender from 2011 to 2020. Panel (a) shows two measures of the gender gap in VC funding: the red line represents the dollar-valued gap in deal size between male- and female-founded startups, while the blue dotted line plots the difference in deal frequency between all-male and women-founded startups (left vertical axis). Panel (b) plots the relative funding gap across VC stages, computed as 1avg. funding for women-foundedavg. funding for all-men $1-\frac{\text{avg. funding for women-founded}}{\text{avg. funding for all-men founded}}$. Lines are regression fits by group. A startup is “women-founded” if at least one founder is a woman but results are qualitatively similar with other classifications.
@@
Collecting detailed data on startups is challenging. Startups are inherently difficult to track compared to publicly listed firms, and, more importantly, there is often limited information on their operations. Who are the founders? What are their gender and background? How much funding have they raised? These challenges complicate productivity comparisons between startups led by female and male founders, as it becomes nearly impossible to control for the quality of a startup. Controlling for both observed and unobserved quality is crucial to identify discrimination in venture capital funding.
@@
We address these problems by studying startups graduating from accelerators. Accelerators can be seen as colleges for startups, as they provide services (e.g., legal support) and mentorship. Similar to college admissions, startups apply to multiple accelerators before enrolling in one of them. We hand-collected information on the universe of startups “graduating” from all U.S. accelerators from their respective websites and other sources (e.g., Linkedin). For each startup, we observe demographic information regarding its founders, funding outcomes over different horizons since graduation, as well as whether they were acquired or went public. We observe various characteristics for accelerators as well, such as their yearly cohort.
Our approach to investigating the gender gap is divided into two steps. In the first step, we build and estimate a one-to-many matching model with non-transferable utility between startups and accelerators. The model estimates the degree of assortative matching, so that high-quality startups and accelerators are more likely to match during the admission process. It is also flexible enough to allow startups to match with accelerators from other states. The second step takes advantage of the breadth of our dataset and estimates outcomes at the startup level, treating the gender of the startup founders and other covariates as factors in a control function approach. This method uses the matching probabilities between accelerators and startups—estimated in the first step—to control for unobserved endogenous quality.
We find strong evidence of a significant gender penalty. As shown in Figure 1, startups with a female founder have a -3.8 percentage point lower probability of reaching the five-million-dollar funding milestone one year after graduation, with the gap widening to -16 percentage points within five years. Since later VC rounds tend to be larger than earlier ones, this widening probability gap translates into an even larger gap in U.S. dollar terms. This result is particularly concerning because attending an accelerator serves as a signal of quality, meaning the quality heterogeneity among the startups in our sample should be smaller than the heterogeneity underlying Figure 1.
The mechanism behind this disparity seems to be related to the age of female founders when attending accelerators. We find that the gender gap is driven by startups where female founders are at an age when they are likely to have young children. These founders face significant challenges in relocating to other states to match with their preferred accelerators, resulting in a lower probability of securing funding. In contrast, female-founded startups that do relocate, while still trailing in early achievements, manage to close most of the gap by the fifth year post-graduation. Network externalities, such as the size of the cohort and the quality of the accelerators, help close the gap. This finding opens new research avenues, suggesting that accelerators not only foster firm growth but may also alleviate the frictions contributing to the gender gap in VC funding. For example, while cohort size may be influenced by the overall structure of the accelerator, access to data on accelerator operations could help researchers understand the services and mentorship they offer, potentially applying these insights to lower-quality accelerators.
Previous research has primarily focused on the probability of securing any investment or smaller amounts, with less attention given to the gender gap in larger VC deals. Understanding this gap is crucial, as biases in larger investments may discourage early-stage investors from funding women-founded startups, perpetuating statistical discrimination (e.g., Ewens and Townsend,, 2020; Gornall and Strebulaev,, 2024; Gafni et al.,, 2021; Zhang,, 2023). Addressing this issue is important not only for ethical reasons, as reducing the gender gap can foster a more equitable environment for women entrepreneurs, but also because doing so generates economic value: despite raising less funding than qualitatively comparable firms, they are no more likely to go bankrupt in the first five years. [^1] Relatedly, a growing literature studies the role of women as directors and exectives on corporate economic and social outcomes (Ferrari et al.,, 2022; Huang and Kisgen,, 2013; Field et al.,, 2020; Adams and Ferreira,, 2009; Kenneth and Dittmar,, 2012). Moreover, they have the same likelihood of being acquired as their male counterparts, potentially indicating a more efficient use of funds. Therefore, closing the gender gap in financing is crucial not only for promoting fairness but also for sustaining innovation and fostering the entry of new firms into highly innovative sectors, such as the technology industry, which is the focus of our study.
Our approach emphasizes the link between accelerator evaluations and VC investment decisions. Investor-led accelerators, such as Y Combinator, are for-profit programs managed by successful VC investors and entrepreneurs (Cohen and Hochberg,, 2014). Funded by venture investors, these accelerators select high-growth potential startups and take equity stakes (Clarysse et al.,, 2016). Graduates from these programs often attract significant VC interest, with accelerator alumni receiving about one-third of Series A funding in the U.S. [^2] See reports by Ian Hathaway. https://www.ianhathaway.org/blog/2019/4/9/accelerated-companies-at-series-a. in 2018. By 2021, Y Combinator’s top 30 graduates were valued at 575 billion US\$, while TechStars and 500 Startups had portfolios worth 71 billion US\$ and 49 companies valued at over 1 billion US\$.
This link is crucial for controlling unobserved startup and accelerator quality. If men-founded startups generally have higher quality during accelerator admission, women-founded startups securing admission to high-quality programs would inherently possess higher, unobservable quality. This unobserved quality can affect startup performance in the VC market, creating endogeneity by correlating gender with the error term in venture performance. To address this, we adopt the modeling approach of Sørensen, (2007), where accelerator assessments correlate with post-accelerator performance. The key assumption is that accelerator outcomes depend on the presence of other participants, while post-program performance is driven by the startup and its associated accelerator.
@@
We estimate the model in two stages. In the first stage, accelerators and startups match based on maximizing expected startup value (e.g., NPV) at graduation. The expected value depends on accelerator characteristics (program quality), startup characteristics, complementary factors (e.g., location), and potentially unobserved qualities. The matching process is competitive and capacity-constrained, with equilibrium determined by pairwise stability, as in (Roth and Sotomayor,, 1990). [^3] This approach differs substantially from the standard discrete choice approach. In a discrete choice model where startups select accelerators, accelerators do not face capacity constraints. As a result, the market competition encountered by startups is not accounted for, which can lead to biased estimates.
In the second stage, fundraising depends on observable variables and an unobserved match quality term, reflecting the tendency of better startups to pair with better accelerators. Drawing from the control function literature (Heckman,, 1974), we show that this unobserved term correlates with observable matching outcomes obtained from the estimation of the matching model. Thus, a contribution of our paper is to account for unobserved heterogeneity in future outcomes related to the matching stage by leveraging the structure of the matching process, an approach that can be exported to study other questions (Akkus et al.,, 2020; Lu and Rui,, 2018).
A common explanation for gender differences is relocation preferences, where women may face location constraints or family obligations that limit their ability to contact investors or fully dedicate themselves to their startups (e.g., Barbanchon et al.,, 2021; Bertrand et al.,, 2010; Zandberg,, 2021; Jayachandran et al.,, 2024). These make female entrepreneurs more likely to locate their businesses in more segregated areas, with shorter commuting times and smaller networks (Rosenthal and Strange,, 2012). Our results point to similar mechanisms in place for accelerator-startup matches, as accelerators with more enrolled startups and thus greater networking opportunities record smaller gender gaps in funding. [^4] Risk aversion is another common explanation of gender differences (Croson and Gneezy,, 2009; Buser et al.,, 2014), but prior research suggests limited evidence linking risk preferences to funding outcomes (Ewens,, 2023), with Gafni et al., (2021) showing that gender differences in risk aversion disappear when accounting for industry factors.
Our results also indicate that the gender gap is unlikely to be explained by differences in pitch styles or networking abilities (Brooks et al.,, 2014; Hu and Ma,, 2021; Exley and Kessler,, 2022). Accelerators assess pitch styles during the admission process (Hallen et al.,, 2020; Stross,, 2012) and provide pitch training (Clingingsmith et al.,, 2022), which mitigates gender-based differences. However, our analysis shows that a gender gap persists in larger investments even after controlling for accelerator quality. Additionally, while female investors may offer better support to women founders (Raina,, 2021; Hebert,, 2023), female founders still face challenges breaking the glass ceiling in accelerators led by women.
Finally, this paper broadly relates to the extensive literature on the gender gap in the labor market (e.g., Biasi and Sarsons,, 2022; Campa et al.,, 2011; Dahl et al.,, 2021; Ghazala and Ferrer,, 2017; Kawaguchi,, 2007) and in the sciences, as most of the founders in our datasets have STEM backgrounds (Iaria et al.,, 2024; Ahn et al.,, 2024; Galasso and Profeta,, 2024). It relates to the literature on the “glass ceiling”—an invisible but persistent barrier that prevent certain groups—particularly women and minorities—from advancing to the highest levels of their professions, despite having the necessary qualifications and abilities (Albrecht et al.,, 2003; Matsa and Miller,, 2011; Bertrand et al.,, 2018).
@@ 2 Data
We define a “program” as a cohort of startups. Some accelerators operate multiple programs across different locations and time periods. In total, we identify 74 programs from 27 accelerators, covering 736 startup graduates.
Table 1 presents summary statistics for the accelerators in our dataset. Around 37% of the programs were located in startup hubs (CA, MA, NY), closely reflecting the broader geographic distribution of U.S. accelerators, with approximately 40% concentrated in major tech hubs such as Silicon Valley, Boston-Cambridge, and New York.
Table 1: Accelerator profiles
Number of Accelerators 27
Number of Programs without Female Founder(s) 56
Number of Programs with Female Founder(s) 18
Number of States Represented 21
Programs in Startup Hubs (CA, NY, MA) 27
Average Cohort Size 17.70
Number of Startup Graduates 736
Table 2 presents summary statistics for accelerator participants. [^6] Industries are classified using the six-digit 2012 NAICS codes: IT Service (519190), Software (511210), Data Processing and Hosting (518210), Internet and Web (519130), and Others (e.g., Healthcare, Mobile Devices). Overall, gender differences are small, with differences across most variables being not statistically significant at the 10% level. However, women-founded startups have fewer founders with Engineering or Science degrees, smaller founding teams, older founders, and are less likely to operate in Software. We control for these differences in our analysis.
@@
All Women Founded Men Founded
Mean St. dev. Mean St. dev. Mean St. dev.
Female founder 0.086 0.280 - - - -
Startup age 0.762 0.899 0.889 1.018 0.750 0.886
At least one serial founder 0.376 0.485 0.286 0.455 0.385 0.487
Founding team size 2.264 0.788 2.016 0.729 2.287 0.790
Team average age 28.776 5.568 30.630 6.651 28.603 5.430
At least one graduate degree 0.353 0.478 0.429 0.499 0.346 0.476
At least one Ph.D. degree 0.075 0.263 0.111 0.317 0.071 0.258
At least one engr/sci degree 0.644 0.479 0.317 0.469 0.675 0.469
By industry:
- IT services 0.382 0.486 0.317 0.469 0.388 0.488
- Software 0.186 0.389 0.222 0.419 0.183 0.387
- Data processing&hosting 0.268 0.443 0.317 0.469 0.263 0.441
- Internet&web 0.092 0.290 0.032 0.177 0.098 0.298
- Others 0.072 0.259 0.111 0.317 0.068 0.253
Observations 736 63 673
  • Notes: This table present summary statistics for the startups in our sample. The term “serial founder” indicates a founder who had already founded at least another startup before the focal one.
@@ 3.1 What are accelerators?
Accelerators, also known as “seed accelerators” or “startup accelerators,” primarily target early-stage, high-tech startups, especially in IT-related industries. During our data period, most accelerators admitted startups ready to raise VC, typically taking a 5% equity stake. [^7] Accelerators also provided seed funding, often considered a stipend for participating entrepreneurs (Cohen and Hochberg,, 2014). See Appendix A.1 for details on accelerator operations.
@@
Graduation Year)
Average Yearly Funding for Cohort: 2008 2009 2010 2011
Panel A: In the first year since graduation
Accelerator graduates 0.82 1.37 1.36 1.45
Comparable startups (Seed and Angel stages) 1.06 1.00 0.86 0.87
Panel B: In the first five years since graduation
Accelerator graduates 1.55 6.70 2.42 2.06
Comparable startups (same age) 2.64 2.05 1.96 1.95
Values in US\$ millions.
  • Notes: This table compares funding received by accelerator graduates with comparable startups in the VC market. Panel A reports the average deal size for accelerator graduates from 2008 to 2011 in their first year post-graduation, alongside startups of similar age in the “seed” or “angel” stages of the VC market for the corresponding year. Panel B presents the average yearly funding over a five-year horizon, calculated as total funding over five years divided by years in operation. The comparison sample consists of startups in the “seed” to “late” stages from PitchBook. Funding amounts are inflation-adjusted to constant dollars.
3.2 Accelerators’ Admission of Women-Founded Startups
This section shows that: (1) there is no evidence that accelerators favored women-founded startups during our data period, (2) accelerators did not admit more women than VCs, and (3) women-founded accelerator participants performed similarly to their male-founded counterparts. These results corroborate our modeling assumptions in the next section.
Accelerators were relatively new in the venture market, with approximately 800 graduates by the end of 2011. Gender diversity was not a selection criterion at the time. For instance, Stross, (2012) notes that Y Combinator focused solely on startup growth potential and did not attempt to balance gender in applications. Criticism for admitting too few women only emerged post-2012. Y Combinator, often scrutinized for its low admission rate of women-founded startups, faced such criticism in 2013.888See: http://www.paulgraham.com/ff.html.
To confirm that public opinion had little influence on accelerator admissions during our data period, we follow Giuli and Kostovetsky, (2014) and regress the likelihood of admitting women-founded startups on the political preferences of the local public during the 2008 presidential election. As shown in Appendix Table A1, we find no significant association between state-level political preferences and the admission of women-founded startups.

### v2

# secReferences in the Online Appendix

@@ Abstract
We examine the growing gender gap in venture capital funding, focusing on accelerator programs in the U.S. We collect a unique dataset with detailed information on accelerators and startups. Using a two-stage methodology, we first estimate a matching model between startups and accelerators, and then use its output to analyze the gender gap in post-graduation outcomes through a control function approach. Our results suggest that female-founded startups face a significant funding disadvantage due to relocation challenges tied to family obligations. However, larger cohorts and higher-quality accelerators help reduce this gap by potentially offering female founders better networking opportunities and mentorship.
@@ 1 Introduction
Notes: Using data from Pitchbook, (2021, 2022), we compute VC deal counts and average deal sizes by gender from 2011 to 2020. Panel (a) shows two measures of the gender gap in VC funding: the red line represents the dollar-valued gap in deal size between male- and female-founded startups, while the blue dotted line plots the difference in deal frequency between all-male and women-founded startups (left vertical axis). Panel (b) plots the relative funding gap across VC stages, computed as $1-\frac{\text{avg. funding for women-founded}}{\text{avg. funding for all-men founded}}$. Lines are regression fits by group. A startup is “women-founded” if at least one founder is a woman but results are qualitatively similar with other classifications.
@@
Collecting detailed data on startups is challenging. Startups are more difficult to track compared to publicly listed firms, and, more importantly, there is often limited information on their operations. Who are the founders? What are their gender and background? How much funding have they raised? These challenges complicate productivity comparisons between startups led by female and male founders, as it becomes nearly impossible to control for the quality of a startup, which is crucial to identify discrimination.
@@
We address these problems by studying startups graduating from accelerators. Accelerators can be seen as “colleges for startups,” as they provide services (e.g., legal support) and mentorship. Similar to college admissions, startups apply to multiple accelerators before enrolling in one of them. We hand-collected information on the universe of startups “graduating” from all U.S. accelerators from their respective websites and other sources such as Linkedin. For each startup, we observe demographic information regarding its founders, funding outcomes over different horizons since graduation, as well as whether they were acquired or went public. We observe various characteristics for accelerators as well, such as their yearly cohort.
Because accelerators selectively admit startups based on potential and because startups apply to accelerators based on fit, the match between a startup and its accelerator contains valuable information about the quality of each enrollment match. In fact, standard regression approaches that ignore this selection process risk conflating founder gender with latent quality differences. To address this, we adopt a two-step empirical strategy that explicitly models the startup–accelerator matching process (e.g., Sørensen, , 2007).
@@
In the first step, we estimate a one-to-many matching model with non-transferable utility between startups and accelerators. This model captures the assortative nature of admissions: high-quality startups tend to match with high-quality accelerators. Importantly, our framework allows startups to match with accelerators outside their home region, enabling us to analyze geographic mobility. The matching process is competitive and capacity-constrained, with equilibrium determined by pairwise stability (Roth and Sotomayor, , 1990). [^2] This approach differs substantially from the standard discrete choice approach. In a discrete choice model where startups select accelerators, accelerators do not face capacity constraints. As a result, the market competition encountered by startups is not accounted for, which can lead to biased estimates. The resulting matching probabilities provide a proxy for unobserved startup quality and the competitive intensity of the admissions process.
In the second step, we analyze post-graduation VC funding outcomes using a control-function approach. By incorporating the estimated match-level information from the first step, we control for endogenous selection into accelerators. This allows us to isolate the effect of founder gender on funding success, net of both observed and unobserved quality differences (Lu and Rui, , 2018; Akkus et al., , 2020).
Our approach introduces a novel two-stage estimation procedure that separates the matching process from the outcome equation. The key advantage of this separation is that it enables a more encompassing definition of markets, obviating the need to segment the U.S. into multiple local markets. This is particularly important in our setting, where 27.17 percent of startups relocate to accelerators outside their home states, rendering localized market definitions impractical. While this comes at the cost of some statistical efficiency compared to joint estimation, it greatly simplifies the likelihood function in the first stage. With such analytical tractability, we do not need computationally intensive simulations, which would require drawing error terms for an astronomical number of possible matches to evaluate the likelihood function. [^3] For instance, a market with 100 startups, 20 accelerators with 5 slots each has more than 10100 possible matches.
The separation of stages also facilitates the application of a control function approach in the second stage, enabling flexible regression-based analysis across various outcome variables. This structure is computationally convenient: by estimating the first stage only once, we can bootstrap the second-stage standard errors without repeatedly solving the matching problem. Although this two-step procedure sacrifices some asymptotic efficiency relative to fully joint estimation methods (e.g., Sørensen, , 2007), it enables scalable inference in the presence of large market sizes.
We find strong evidence of a significant gender penalty. Startups with a female founder have a -3.8 percentage point lower probability of reaching the five-million-dollar funding milestone one year after graduation, with the gap widening to -16 percentage points within five years. Since later VC rounds tend to be larger than earlier ones, this widening probability gap translates into an even larger gap in U.S. dollar terms. This result is particularly concerning because attending an accelerator serves as a signal of quality, meaning the quality heterogeneity among the startups in our sample should be smaller than the heterogeneity underlying Figure 1, which draws from the universe of U.S. startups.
What mechanisms drive this post-accelerator gender gap? Our analysis points to geographic mobility constraints as a key factor. The disadvantage for female founders is most pronounced among those in their late 20s to early 30s, an age range when many face family obligations or childcare responsibilities. These constraints make relocation more difficult, leading some women to enroll in less prominent local accelerators or to forgo the opportunity altogether. In turn, this limits their access to investor networks and follow-on capital.
At the same time, we find that female-founded startups that do relocate, overcoming family-related location frictions, continue to lag male-founded startups in early fundraising, but nearly close the gap within five years after graduation. This convergence is especially strong when women-led ventures participate in accelerators with large cohorts or top-tier reputations, which likely offer richer mentorship and broader investor exposure. These results suggest that accelerator programs can partially mitigate the gender funding gap: by providing peer support and access to professional networks, they help female entrepreneurs overcome disadvantages linked to mobility constraints and narrower informal connections.
Our findings carry important implications. First, we contribute to research on the mechanisms behind gender disparities in entrepreneurship. While prior work has emphasized behavioral differences (Croson and Gneezy, , 2009), statistical discrimination, or gendered pitch perceptions (Kanze et al., , 2018; Exley and Kessler, , 2022), our findings highlight geographic mobility constraints—often tied to family responsibilities—as an underappreciated barrier for female founders. This complements evidence from labor and urban economics on how commuting costs and caregiving roles shape women’s labor market choices and entrepreneurial activity (Bertrand et al., , 2010; Rosenthal and Strange, , 2012; Le Barbanchon et al., , 2021; Jayachandran, , 2021; Zandberg, , 2021). Crucially, we show that even when women overcome these frictions and access top-tier accelerator programs, funding gaps persist —suggesting that structural constraints, rather than pitch style or ambition, are central to the problem. Our analysis also connects to broader work on the “glass ceiling” in both labor markets and science, where similarly qualified women face persistent barriers to advancement (Bertrand et al., , 2018; Galasso and Profeta, , 2024). [^4] This paper also broadly relates to the extensive literature on the gender gap in the labor market (e.g., Biasi and Sarsons, , 2022; Campa et al., , 2011; Dahl et al., , 2021; Ghazala and Ferrer, , 2017; Kawaguchi, , 2007) and in the sciences, as most of the founders in our datasets have STEM backgrounds (Iaria et al., , 2024; Ahn et al., , 2024; Galasso and Profeta, , 2024).
Second, we show that certain features of the entrepreneurial ecosystem can alleviate this disparity. In our data, female-led ventures are just as likely to survive and to be acquired as their male-led counterparts, despite receiving less early-stage funding (e.g., Ewens and Townsend, , 2020; Gornall and Strebulaev, , 2024). This suggests that capital constraints on women-founded startups represent lost growth opportunities rather than rational responses to poorer performance (Huang and Kisgen, , 2013; Nanda and Rhodes-Kropf, , 2013). Closing the gender gap in financing, therefore, is not only a matter of equity but also one of economic efficiency: high-potential ventures led by women may be underfunded relative to their merits, which hampers innovation and limits new firm entry. More broadly, such underfunding may reduce the variety of business ideas brought to market (Pistilli et al., , 2023).
Finally, our study sheds light on how organizational interventions can promote inclusion. Accelerators—especially investor-driven programs like Y Combinator or TechStars—have become a critical gateway to venture funding (Hallen et al., , 2020; Yu, , 2020). Our results suggest that these programs, through their cohort-based mentorship and investor networks, serve as an equalizing force that can partially counteract gender-based network gaps (Clingingsmith et al., , 2022; Cohen et al., , 2019). This opens up new avenues for research and practice: for example, accelerator managers and policy makers might consider cohort policies or support services (such as child-care assistance or remote program options) to encourage greater participation of female founders who face mobility constraints (Balachandra, , 2019; Brush and Elam, , 2024). In sum, by identifying relocation frictions as a barrier and accelerators as a tool to overcome it, our study contributes a deeper understanding of why female entrepreneurs remain underfunded and what can be done to bridge this gap.
@@ 2 Data
We define a “program” as a cohort of startups. Some accelerators operate multiple programs across different locations and time periods. In total, we identify 74 programs from 27 accelerators, covering 736 startup graduates (the average cohort size is 17.7 startups). 18 of these programs have startups with at least a female founder. Around 37% of the programs were located in startup hubs (CA, MA, NY), closely reflecting the broader geographic distribution of U.S. accelerators, with approximately 40% concentrated in major tech hubs such as Silicon Valley, Boston-Cambridge, and New York.
Table 1 presents summary statistics for accelerator participants. [^6] Industries are classified using the six-digit 2012 NAICS codes: IT Service (519190), Software (511210), Data Processing and Hosting (518210), Internet and Web (519130), and Others (e.g., Healthcare, Mobile Devices). Overall, gender differences are small, with differences across most variables being not statistically significant at the 10% level. However, women-founded startups, meaning startups with at least one female founder, have fewer founders with Engineering or Science degrees, smaller founding teams, older founders, and are less likely to operate in Software. We control for these differences in our analysis.
@@
All Women Founded Men Founded Founding Team Startup age (years) 0.762 0.889 0.750 At least one serial founder 0.376 0.286 0.385 Team size 2.264 2.016 2.287 Average age of founders 28.776 30.630 28.603 At least one graduate degree 0.353 0.429 0.346 At least one Ph.D. 0.075 0.111 0.071 At least one engineering/science degree 0.644 0.317 0.675 Industry IT services 0.382 0.317 0.388 Software 0.186 0.222 0.183 Data processing & hosting 0.268 0.317 0.263 Internet & web 0.092 0.032 0.098 Other industries 0.072 0.111 0.068 Observations 736 63 673 Notes: This table presents summary statistics for the startups in our sample. “Women Founded” refers to startups with at least one female founder. “Men Founded” are startups with all-male founding teams. “Serial founder” refers to a founder who had previously created at least one other startup.
@@ 3.1 What are accelerators?
Accelerators, also known as “seed accelerators” or “startup accelerators,” primarily target early-stage, high-tech startups, especially in IT-related industries. During our data period, most accelerators admitted startups ready to raise VC, typically taking a 5% equity stake.See Appendix A.1 for details on accelerator operations.
@@
Graduation Year Average Yearly Funding for Cohort: 2008 2009 2010 2011 Panel A: In the first year since graduation Accelerator graduates 0.82 1.37 1.36 1.45 Comparable startups (Seed and Angel stages) 1.06 1.00 0.86 0.87 Panel B: In the first five years since graduation Accelerator graduates 1.55 6.70 2.42 2.06 Comparable startups (same age) 2.64 2.05 1.96 1.95 Values in US\$ millions. Notes: This table compares funding received by accelerator graduates with comparable startups in the VC market. Panel A reports the average deal size for accelerator graduates from 2008 to 2011 in their first year post-graduation, alongside startups of similar age in the “seed” or “angel” stages of the VC market for the corresponding year. Panel B presents the average yearly funding over a five-year horizon, calculated as total funding over five years divided by years in operation. The comparison sample consists of startups in the “seed” to “late” stages from PitchBook. Funding amounts are inflation-adjusted to constant dollars.
3.2 Accelerators’ Admission and Gender Diversity
Accelerators were relatively new in the venture market, with approximately 800 graduates by the end of 2011. Gender diversity was not a selection criterion at the time. For instance, Stross, (2012) notes that Y Combinator focused solely on startup growth potential and did not attempt to balance gender in applications. Criticism for admitting too few women only emerged post-2012. Y Combinator, often scrutinized for its low admission rate of women-founded startups, faced such criticism in 2013.777See: http://www.paulgraham.com/ff.html.
To test whether public opinion influenced accelerator admissions over our sample period, we regress the probability of admitting a women‐founded startup on state‐level vote share for the Democratic presidential candidate in 2008, following Giuli and Kostovetsky, (2014). The hypothesis is that accelerators located in states with stronger Democratic support would be more sensitive to gender bias. As shown in Appendix Table A1, we find no statistically significant relationship between Democratic vote share and the likelihood of admitting women‐founded startups, suggesting that there is no evidence that accelerators favored women founders during our sample period.
Therefore, we rule out the possibility that accelerators showed systematically greater favoritism toward certain types of startups—such as those founded by women—in a way that would alter the interpretation and external validity of our results with respect to the gender gap in financing.